# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ Appendix A Proofs for Section 3
$\begin{split}\min_{x\in\mathbb{R}^{6}}&~{}~{}\frac{1}{2}\sum_{i=1}^{6}x_{i}^{2}+\sum_{i=1}^{6}x_{i},\\ \text{s.t.}&~{}~{}x_{1}+x_{2}\geq 1,\ x_{2}+x_{3}\geq 1,\ x_{3}+x_{4}\geq 1,\\ &~{}~{}x_{4}+x_{5}\geq 1,\ x_{5}+x_{6}\geq 1,\ x_{6}+x_{1}\geq 1,\\ &~{}~{}x_{j}\in\{0,1\},\ \forall~{}j\in\{1,2,\dots,6\}.\end{split}$ $\begin{split}\min_{x\in\mathbb{R}^{6}}&~{}~{}\frac{1}{2}\sum_{i=1}^{6}x_{i}^{2}+\sum_{i=1}^{6}x_{i},\\ \text{s.t.}&~{}~{}x_{1}+x_{2}\geq 1,\ x_{2}+x_{3}\geq 1,\ x_{3}+x_{1}\geq 1,\\ &~{}~{}x_{4}+x_{5}\geq 1,\ x_{5}+x_{6}\geq 1,\ x_{6}+x_{4}\geq 1,\\ &~{}~{}x_{j}\in\{0,1\},\ \forall~{}j\in\{1,2,\dots,6\}.\\ ~{}\end{split}$ $\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$
$\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$
Denote $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ as the graph representations of the above two MI-LCQP problems. Let $s_{i}^{l},t_{j}^{l}$ and $\hat{s}_{i}^{l},\hat{t}_{j}^{l}$ be the features at the $l$-th layer when apply a GNN $F\in\mathcal{F}_{\textup{MI-LCQP}}$ to $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$. We will prove by induction that for any $0\leq l\leq L$, the followings hold:
@@
It is clear that the conditions (a) and (b) are true for $l=0$, since $v_{i}=\hat{v}_{i}$ is constant in $i\in\{1,2,\dots,6\}$, and $w_{j}=\hat{w}_{j}$ is constant in $j\in\{1,2,\dots,6\}$. Now suppose that the conditions (a) and (b) are true for $l-1$ where $1\leq l\leq L$. We denote that $s^{l-1}=s_{i}^{l-1}=\bar{s}_{i}^{l-1},\ \forall~{}i\in\{1,2,\dots,6\}$ and $t^{l-1}=t_{j}^{l-1}=\hat{t}_{j}^{l-1},\ \forall~{}j\in\{1,2,\dots,6\}$. It can be computed for any $i\in\{1,2,\dots,6\}$ and $j\in\{1,2,\dots,6\}$ that
@@
$\begin{split}\min_{x\in\mathbb{R}^{7}}&~{}~{}\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~{}~{}x_{1}-x_{2}=0,\ x_{2}-x_{1}=0,\\ &~{}~{}x_{3}-x_{4}=0,\ x_{4}-x_{5}=0,\\ &~{}~{}x_{5}-x_{6}=0,\ x_{6}-x_{7}=0,\ x_{7}-x_{3}=0,\\ &~{}~{}x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~{}~{}0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~{}j\in\{1,2,\dots,7\}.\end{split}$ $\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
@@
$\begin{split}\min_{x\in\mathbb{R}^{7}}&~{}~{}\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~{}~{}x_{1}-x_{2}=0,\ x_{2}-x_{3}=0,\ x_{3}-x_{1}=0,\\ &~{}~{}x_{4}-x_{5}=0,\ x_{5}-x_{6}=0,\\ &~{}~{}x_{6}-x_{7}=0,\ x_{7}-x_{4}=0,\\ &~{}~{}x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~{}~{}0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~{}j\in\{1,2,\dots,7\}.\end{split}$ $\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
One can easily verify that both problems are feasible. The feasible region of the first problem is $\{(3,3,0,0,0,0,0)\}$ and the feasible region of the second problem is $\{(2,2,2,0,0,0,0)\}$. Thus, the two problems have the same optimal objective value, say $24$, but their optimal solution sets are disjoint, without any common element.
@@
Appendix C Characterization of unfoldability and MP-tractability
In this section, we discuss some further characterizations of the unfoldability and the MP-tractability for MI-LCQP-graphs defined in Section 4.2.
We first prove that the unfoldability implies the MP-tractability but they are not equivalent.
Proposition C.1.
If $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ is unfoldable, then it is also MP-tractable.
Proof.
Let $(\mathcal{I},\mathcal{J})$ be the final stable partition of $V\cup W$ generated by WL test on $G_{\textup{MI-LCQP}}$ without collision, where =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ is a partition of $V=\{1,2,\dots,m\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ is a partition of $W=\{1,2,\dots,n\}$. Since we assume that $G_{\textup{MI-LCQP}}$ is foldable, we have $t=n$ and $|J_{1}|=|J_{2}|=\cdots=|J_{n}|=1$. Then for any $q,q^{\prime}\in\{1,2,\dots,t\}$, the submatrix $(Q_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$ is a $1\times 1$ matrix and hence has identical entries.
$C_{i}^{L,V}=C_{i^{\prime}}^{L,V}$
which implies that the submatrix $(A_{ij})_{i\in I_{p},j\in J_{q}}$ has identical entries since $|J_{q}|=1$. Therefore, $G_{\textup{MI-LCQP}}$ is MP-tractable. ∎
Proposition C.2.
There exist MP-tractable instances in $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ that are not unfoldable.
Proof.
$\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
[…]
$\min$ $\frac{1}{2}x_{2}^{2}+x_{1}+x_{2}+x_{3}$
s.t. $x_{1}+x_{3}\leq 1$
$x_{1}-x_{2}+x_{3}\leq 1$
$0\leq x_{1},x_{2},x_{3}\leq 1$
$x_{1},x_{2},x_{3}\in\mathbb{Z}$
MI-LCQP problem
Figure 6. Example for proving Proposition C.2
Consider the example in Figure 6, for which the final stable partition is =$\mathcal{I}=\{\{1\},\{2\}\}$ and $\mathcal{J}=\{\{1,3\},\{2\}\}$. It is not unfoldable since the class $\{1,3\}$ in $\mathcal{J}$ has two elements. However, it is MP-tractable since $A_{11}=A_{13}=1$ and $A_{21}=A_{23}=1$. ∎
We then show that a generic MI-LCQP-graph in $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ must be unfoldable. Intuitively, if $c\in\mathbb{R}^{n}$ is randomly sampled from a continuous distribution with density, then almost surely it holds that $x_{j}\neq x_{j^{\prime}}$ for any $j\neq j^{\prime}$, which implies that the vertices in $W$ have different colors initially and always, if there are no collisions of hash functions.
Proposition C.3.
Let $\mathbb{P}$ be a probability measure over $\mathcal{G}_{\textup{MI-LCQP}}$ such that the marginal distribution $\mathbb{P}_{c}$ of $c\in\mathbb{R}^{n}$ has density. Then $\mathbb{P}[\mathcal{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{unfold}}^{m,n}]=1$.
Proof.
where we assumed that no collisions happen in hash functions. Therefore, any $j,j^{\prime}\in W$ with $j\neq j^{\prime}$ are not the in same class of the final stable partition $(\mathcal{I},\mathcal{J})$, which proves the unfoldability. ∎
As a direct corollary of Proposition C.1 and Proposition C.3, a generic MI-LCQP-graph in $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ must also be MP-tractable.
Corollary C.4.
Let $\mathbb{P}$ be a probability measure over $\mathcal{G}_{\textup{MI-LCQP}}$ such that the marginal distribution $\mathbb{P}_{c}$ of $c\in\mathbb{R}^{n}$ has density. Then $\mathbb{P}[\mathcal{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MP}}^{m,n}]=1$.
Appendix D Proofs for Section 4.2
@@ Appendix C Characterization of unfoldability and MP-tractability
This section collects the proofs of Theorems 4.6, 4.7, and 4.9. Similar to the LCQP case, the proofs are also based on the WL test (Algorithm 1) and its separation power to distinguish MI-LCQP problems with different properties. We define the separation power of Algorithm 1 as follows.
Definition D.1.
@@
Let $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ be two MI-LCQP-graphs and denote $\{\{C_{i}^{L,V}\}\}_{i=0}^{m},\{\{C_{j}^{L,W}\}\}_{j=0}^{n}$ and $\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m},\{\{\hat{C}_{j}^{L,W}\}\}_{j=0}^{n}$ as the color multisets output by Algorithm 1 on $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$.
  1. (a)
    We say $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$ if for all $L\in\mathbb{N}$ and all hash functions, $\{\{C_{i}^{L,V}\}\}_{i=0}^{m}=\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m}$ and $\{\{C_{j}^{L,W}\}\}_{j=0}^{n}=\{\{\hat{C}_{j}^{L,W}\}\}_{j=0}^{n}$.
  2. (b)
@@
    We say $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$ if for all $L\in\mathbb{N}$ and all hash functions, $\{\{C_{i}^{L,V}\}\}_{i=0}^{m}=\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m}$ and $C_{j}^{L,W}=\hat{C}_{j}^{L,W},\ \forall~{}j\in\{1,2,\dots,n\}$.

### v2

@@ Appendix A Proofs for Section 3
[…]
Denote $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$ as the graph representations of the above two MI-LCQP problems. Let $s_{i}^{l},t_{j}^{l}$ and $\hat{s}_{i}^{l},\hat{t}_{j}^{l}$ be the attributes at the $l$-th layer when apply a GNN $F\in\mathcal{F}_{\textup{MI-LCQP}}$ to $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$. We will prove by induction that for any $0\leq l\leq L$, the followings hold:
@@
It is clear that the conditions (a) and (b) are true for $l=0$, since $v_{i}=\hat{v}_{i}=(1,\geq)$ is constant in $i\in\{1,2,\dots,6\}$, and $w_{j}=\hat{w}_{j}=(1,0,1,1)$ is constant in $j\in\{1,2,\dots,6\}$. Now suppose that the conditions (a) and (b) are true for l1l-1 where $1\leq l\leq L$. We denote that $s^{l-1}=s_{i}^{l-1}=\bar{s}_{i}^{l-1},\ \forall~i\in\{1,2,\dots,6\}$ and $t^{l-1}=t_{j}^{l-1}=\hat{t}_{j}^{l-1},\ \forall~j\in\{1,2,\dots,6\}$. It can be computed for any $i\in\{1,2,\dots,6\}$ and $j\in\{1,2,\dots,6\}$ that
@@
We also remark that, although the two MI-LCQP graphs with distinct optimal objective values presented above are indistinguishable by GNNs considered in this paper, they can be distinguished by $3$-hop GNNs [feng2022powerful, chen2025expressive]. Crucially, when we scale these examples to $10$ variables and $10$ constraints (one graph with 20 connected nodes and the other with two 10-node components), $3$-hop GNNs fail while $5$-hop GNNs succeed, confirming that larger $k$ in $k$-hop GNNs reduces GNN-unfriendly cases, though they do not solve all MI-LCQPs.
@@
$\begin{split}\min_{x\in\mathbb{R}^{7}}&~~\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~~x_{1}-x_{2}=0,\ x_{2}-x_{1}=0,\\ &~~x_{3}-x_{4}=0,\ x_{4}-x_{5}=0,\\ &~~x_{5}-x_{6}=0,\ x_{6}-x_{7}=0,\ x_{7}-x_{3}=0,\\ &~~x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~~0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~j\in\{1,2,\dots,7\}.\end{split}$ $v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
@@
$\begin{split}\min_{x\in\mathbb{R}^{7}}&~~\frac{1}{2}x^{\top}\mathbf{1}\mathbf{1}^{\top}x+\mathbf{1}^{\top}x,\\ \text{s.t.}&~~x_{1}-x_{2}=0,\ x_{2}-x_{3}=0,\ x_{3}-x_{1}=0,\\ &~~x_{4}-x_{5}=0,\ x_{5}-x_{6}=0,\\ &~~x_{6}-x_{7}=0,\ x_{7}-x_{4}=0,\\ &~~x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=6\\ &~~0\leq x_{j}\leq 3,\ x_{j}\in\mathbb{Z},\ \forall~j\in\{1,2,\dots,7\}.\end{split}$ $v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$$w_{7}$
As we mentioned in Section 4.1, both problems are feasible with the same optimal objective value, but have disjoint optimal solution sets.
@@
Appendix C Proofs for Section 4.2
@@
This section collects the proofs of Theorems 4.6 and 4.7. Similar to the LCQP case, the proofs are also based on the WL test (Algorithm 1) and its separation power to distinguish MI-LCQP problems with different properties. We define the separation power of Algorithm 1 as follows.
Definition C.1.
@@
Let $G_{\textup{MI-LCQP}},\hat{G}_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ be two MI-LCQP-graphs. Denote by $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ and $\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{\hat{C}_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ the color multisets output by Algorithm 1 on $G_{\textup{MI-LCQP}}$ and $\hat{G}_{\textup{MI-LCQP}}$. We say:
  1. (1)
    $G_{\textup{MI-LCQP}}\sim\hat{G}_{\textup{MI-LCQP}}$ if $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m}=\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m}$ and $\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}=\{\!\!\{\hat{C}_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ hold for all $L\in\mathbb{N}$ and all hash functions;
  2. (2)
@@
    $G_{\textup{MI-LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{MI-LCQP}}$ if $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m}=\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m}$ and $C_{j}^{L,W}=\hat{C}_{j}^{L,W},\ \forall~j\in\{1,2,\dots,n\}$, for all $L\in\mathbb{N}$ and all hash functions.